LOIRP Status Report 23 August 2013

Dennis Wingo:
— Status morning of August 23, 2013.
Yesterday I was rummaging through all of our heads and we found one that has some significant life left on it. We went through a long period where we had an anomalous failure that we could never explain where the machine would work just fine but the demodulator would be off frequency, thus preventing the reconstruction of an image. When we would change a head, most of the time it would fix it. Thus we had heads with some life left on them but they would not work.
One of Ken’s friends (Terry) who used to work on FR-900’s told us what was going on and since then we have been able to use some of these half or quarter or three quarter worn heads. We are putting another head on today and hopefully it will have enough life on it to get some more images.
We do this as the head that is mostly dead has had problems on two tapes out of six, and that is anomalously high, so maybe this will solve those problems…..
— Status, end of Day August 23, 2013, 47 years after LO-1-101H
Ran 2.5 tapes today. We had a problem with G5-004 and G5-006 yesterday that we have confirmed is not head related. We can capture these images from Madrid tapes or the other end of Goldstone tapes later.
We have captured G5-007 and G5-010 today. G5-008 and 009 are duplicates of W5-003 and 004. This is what we expected to start seeing but it is gratifying anyway.
G5-007, partial capture LOV-013H, 011M
G5-010, partial capture LOV-019M, complete capture 020H
Next week we will start running more tapes in earnest since we have a head with more meat on it installed.